# Characterizing Multi-Hunk Patches: Divergence, Proximity, and LLM Repair Challenges

### v1

@@ III-E Impact of Augmented Prompting (RQ2)
TABLE VI: Hunk divergence for fixed and unfixed bugs, and spatial proximity distribution (fixed) across model families
Model Hunk Divergence (Fixed) Hunk Divergence (Unfixed) Spatial Proximity (Fixed)
@@
We further examine feedback-based refinement, which augments prompting with dynamic execution information derived from compilation failures or test failures, following prior work [65, 66, 67]. To limit cost, the feedback prompt uses the method-level context and the best retrieval setup (token+text-embed). Despite this conservative design, feedback yields the highest accuracy across all settings. o4-mini achieves 35.75% Plausible@1 and LLaMA3.3 achieves 24.46%, both surpassing their respective static prompt baselines. Notably, the accuracy of o4-mini with feedback surpasses that of the static prompt using file-level context ($28.30\%$), while costing less than half as much (\$14.63 vs. \$43.35). These results indicate that feedback, driven by runtime information from failed tests or execution traces, can guide the model more effectively than simply expanding the prompt context. Moreover, feedback mitigates the limitations of narrow context by incorporating dynamic execution traces, which provide concrete, failure-specific guidance that static prompts lack.
@@ III-F Repair Outcomes by Hunk Divergence and Proximity (RQ3)
TABLE VII: Hunk divergence for fixed and unfixed bugs, and spatial proximity (fixed) across best prompt strategies
Model Technique Variant Divergence Score (Fixed) Divergence Score (Unfixed) Spatial Proximity (Fixed)
@@
Successful repairs are consistently associated with low divergence across all five models (mean 0.26absent0.26$\approx$ 0.26≈ 0.260.280.280.280.28; median 0.24absent0.24$\approx$ 0.24≈ 0.240.290.290.290.29). Variance remains narrow (standard deviation $<0.20$), indicating high semantic alignment across generated patches. In contrast, unfixed bugs exhibit substantially higher hunk divergence (mean: 0.490.490.490.490.530.530.530.53; median: 0.450.450.450.450.530.530.530.53), along with broader variance and higher maxima (up to $1.60$). These results reflect the increased difficulty LLMs face in coordinating complex, semantically diverse edits across scattered code regions.
Table VII reports our analysis across the best-performing prompting configurations, i.e., (1) Token + text-embed, (2) File Context, and (3) Feedback Loop, applied to o4-mini and Llama3.3. Across all strategies, fixed bugs consistently exhibit lower divergence than unfixed bugs. The mean divergence for fixed bugs ranges from $0.28$ to $0.40$ (median: 0.280.280.280.280.390.390.390.39), while unfixed bugs display substantially higher scores (mean: 0.490.490.490.490.560.560.560.56; median: 0.450.450.450.450.530.530.530.53), indicating that greater heterogeneity within hunks correlates with repair failure. Among the evaluated strategies, Feedback Loop achieves the lowest divergence values for fixed bugs in both models (mean: 0.290.290.290.290.300.300.300.30; median: 0.300.300.300.300.310.310.310.31), and the highest percentage of repairs in the Nucleus and Cluster classes. This pattern suggests improved effectiveness on bugs involving spatially localized edits. These results demonstrate that prompting strategy significantly influences both semantic coherence and spatial locality of successful patches. Notably, the Feedback Loop configuration demonstrates the highest effectiveness in achieving low-divergence repairs within high-proximity contexts, thereby facilitating more precise and coordinated multi-hunk program repair.
@@ IV Threats to Validity
Hunk4J is derived exclusively from Java projects, which may limit generalizability to other languages. Although these bugs are real-world and diverse, their distribution reflects the design and testing philosophy of Java projects, which may bias the scope of repair challenges.
@@
Our evaluation assumes perfect fault localization, following prior LLM-based repair studies to isolate patch generation quality [35, 41, 33, 38, 34, 17, 10, 18]. While this sets an upper bound on accuracy, real-world scenarios often involve imperfect or incomplete localization. Notably, even with perfect localization, LLMs struggle with multi-hunk repairs, failing on all Fragment-class bugs (Table VI), highlighting challenges in semantic coordination beyond localization.
@@
To support reproducibility, we release Birch, all scripts, prompts, and the Hunk4J dataset [68].
@@ V Related Work
Most APR benchmarks offer limited support for multi-hunk bugs. While Defects4J [51], Bugs.jar [69], BugsInPy [70], and LMDefects [41] include them, others offer limited coverage [71, 72, 73] or focus on single-hunk edits [74, 75, 76]. Existing multi-hunk LLM repair studies rely on synthetic [41], function-level [42], or algorithmic [40] benchmarks, limiting generalizability. We introduce Hunk4J, a real-world, multi-hunk benchmark derived from Defects4J, augmented with divergence metrics for evaluating LLM-based repair.
Prior studies have applied change metrics for fault prediction. Graves et al. [77] used coarse-grained measures such as file and modification counts; D’Ambros et al. [78] highlighted the predictive power of change coupling and scatter; and Ferzund et al. [79] proposed hunk-level metrics. In contrast, we quantify semantic and structural heterogeneity in multi-hunk patches to assess LLM-based repair effectiveness.
Prior work on multi-hunk repair with LLMs [41, 42, 40, 38, 39, 80, 35] exhibits wide variation in datasets, prompting strategies, and evaluation protocols, often reporting aggregate results without isolating multi-hunk accuracy. While prior studies have examined patch properties such as hunk divisibility [80], they do not account for spatial dispersion or intra-patch heterogeneity. We address this limitation by introducing Hunk Divergence and Spatial Proximity, two metrics that capture lexical, structural, and hierarchical relationships between hunks to enable complexity-aware analysis.

### v2

@@ III-E Impact of Augmented Prompting (RQ2)
TABLE VI: Hunk divergence for fixed and unfixed bugs, and spatial proximity distribution (% fixed) across model families
Model Hunk Divergence (Fixed) Hunk Divergence (Unfixed) Spatial Proximity (% Fixed)
@@
We further examine feedback-based refinement, which augments prompting with dynamic execution information derived from compilation failures or test failures, following prior work [flakidock:icse25, ruan:specrover:icse25, bug-reprodcution]. To limit cost, the feedback prompt uses the method-level context and the best retrieval setup (token+text-embed). Despite this conservative design, feedback yields the highest accuracy across all settings. o4-mini achieves 35.75% Plausible@1 and LLaMA3.3 achieves 24.46%, both surpassing their respective static prompt baselines. Notably, the accuracy of o4-mini with feedback surpasses that of the static prompt using file-level context ($28.30\%$), while costing less than half as much (\$14.63 vs. \$43.35). These results indicate that feedback, driven by runtime information from failed tests or execution traces, can guide the model more effectively than simply expanding the prompt context. Moreover, feedback mitigates the limitations of narrow context by incorporating dynamic execution traces, which provide concrete, failure-specific guidance that static prompts lack.
@@ III-F Repair Outcomes by Hunk Divergence and Proximity (RQ3)
TABLE VII: Hunk divergence for fixed and unfixed bugs, and spatial proximity (% fixed) across best prompt strategies
Model Technique Variant Divergence Score (Fixed) Divergence Score (Unfixed) Spatial Proximity (% Fixed)
@@
Successful repairs are consistently associated with low divergence across all five models (mean $\approx 0.26$$0.28$; median $\approx 0.24$$0.29$). Variance remains narrow (standard deviation <0.20<0.20), indicating high semantic alignment across generated patches. In contrast, unfixed bugs exhibit substantially higher hunk divergence (mean: 0.490.490.530.53; median: 0.450.450.530.53), along with broader variance and higher maxima (up to $1.60$). These results reflect the increased difficulty LLMs face in coordinating complex, semantically diverse edits across scattered code regions.
Table VII reports our analysis across the best-performing prompting configurations, i.e., (1) Token + text-embed, (2) File Context, and (3) Feedback Loop, applied to o4-mini and Llama3.3. Across all strategies, fixed bugs consistently exhibit lower divergence than unfixed bugs. The mean divergence for fixed bugs ranges from $0.28$ to $0.40$ (median: 0.280.280.390.39), while unfixed bugs display substantially higher scores (mean: 0.490.490.560.56; median: 0.450.450.530.53), indicating that greater heterogeneity within hunks correlates with repair failure. Among the evaluated strategies, Feedback Loop achieves the lowest divergence values for fixed bugs in both models (mean: 0.290.290.300.30; median: 0.300.300.310.31), and the highest percentage of repairs in the Nucleus and Cluster classes. This pattern suggests improved effectiveness on bugs involving spatially localized edits. These results demonstrate that prompting strategy significantly influences both semantic coherence and spatial locality of successful patches. Notably, the Feedback Loop configuration demonstrates the highest effectiveness in achieving low-divergence repairs within high-proximity contexts, thereby facilitating more precise and coordinated multi-hunk program repair.
@@ IV Threats to Validity
Hunk4J is derived exclusively from Java projects, which may limit generalizability to other languages. Although these bugs are real-world and diverse, their distribution reflects the design and testing philosophy of Java projects, which may bias the scope of repair challenges. We evaluate on Java because it is a popular language and use the widely adopted Defects4J benchmark. Our metrics, hunk divergence and spatial proximity, and our evaluation steps do not rely on language-specific features, so the underlying insights should apply to other languages; nevertheless, additional experiments on non-Java datasets are required to empirically validate this hypothesis and assess generalizability beyond Java and Defects4J.
Our goal is to study how the structure of real developer-written patches affects LLM repair. We use the original patches from Defects4J because they reflect how developers fixed the bugs. A simpler or semantically equivalent patch may exist for some bugs; however, we intentionally compute hunk divergence and spatial proximity from the actual developer patch, not hypothetical alternatives. This provides us with a consistent and realistic approach to studying what makes multi-hunk repair challenging.
@@
Our evaluation assumes perfect fault localization, following prior LLM-based repair studies to isolate patch generation quality [xia:alpharepair:fse22, fan:code-repair-with-llm:icse23, iter, repair-agent, cedar, glance, reptory, katana]. While this sets an upper bound on accuracy, real-world scenarios often involve imperfect or incomplete localization. Notably, even with perfect localization, LLMs struggle with multi-hunk repairs, failing on all Fragment-class bugs (Table VI), highlighting challenges in semantic coordination beyond localization.
@@
To support reproducibility, we release Birch, all scripts, prompts, and the Hunk4J dataset [anon-repo].
@@ V Related Work
Most APR benchmarks offer limited support for multi-hunk bugs. While Defects4J [defects4j], Bugs.jar [bugsjar:msr18], BugsInPy [bugsinpy:fse20], and LMDefects [fan:code-repair-with-llm:icse23] include them, others offer limited coverage [bears:saner19, manybugs:tse15, bugsjs] or focus on single-hunk edits [chen:codrep:arxiv18, lin:quixbugs:splash17, karampatsis:manysstubs4j:msr20]. Existing multi-hunk LLM repair studies rely on synthetic [fan:code-repair-with-llm:icse23], function-level [srepair:arxiv24], or algorithmic [yang:fine-tuning-for-repair-with-llm:arxiv24] benchmarks, limiting generalizability. We introduce Hunk4J, a real-world, multi-hunk benchmark derived from Defects4J, augmented with divergence metrics for evaluating LLM-based repair.
Prior studies have applied change metrics for fault prediction. Graves et al. [graves2000predicting] used coarse-grained measures such as file and modification counts; D’Ambros et al. [dambros2012evaluating] highlighted the predictive power of change coupling and scatter; and Ferzund et al. [ferzund2009software] proposed hunk-level metrics. In contrast, we quantify semantic and structural heterogeneity in multi-hunk patches to assess LLM-based repair effectiveness.
Prior work on multi-hunk repair with LLMs [fan:code-repair-with-llm:icse23, srepair:arxiv24, yang:fine-tuning-for-repair-with-llm:arxiv24, repair-agent, huang:empirical-fine-tuning-for-apr:ase23, xin:empirical-study-hunk:fse24, xia:alpharepair:fse22] exhibits wide variation in datasets, prompting strategies, and evaluation protocols, often reporting aggregate results without isolating multi-hunk accuracy. While prior studies have examined patch properties such as hunk divisibility [xin:empirical-study-hunk:fse24], they do not account for spatial dispersion or intra-patch heterogeneity. We address this limitation by introducing Hunk Divergence and Spatial Proximity, two metrics that capture lexical, structural, and hierarchical relationships between hunks to enable complexity-aware analysis.